OpenStudy (anonymous):

A=0 19 0 1 what is A^-1

OpenStudy (aa123):

hi

OpenStudy (anonymous):

hi. @aa123

OpenStudy (aa123):

Hmmm

OpenStudy (aa123):

umm what grade math is this

OpenStudy (aa123):

sorry i don't think i know this stuff

OpenStudy (anonymous):

11th, algebra 2

OpenStudy (unklerhaukus):

\[\mathbf {AI}= \left[\begin{matrix}0&19\\0&1\end{matrix}\right|\left.\begin{matrix}1&0\\0&1\end{matrix}\right]\] row reduce this to \[\mathbf {IA}^{-1}= \left[\begin{matrix}1&0\\0&1\end{matrix}\right|\left.\begin{matrix}?&?\\?&?\end{matrix}\right]\]

OpenStudy (aa123):

@UnkleRhaukus hi

OpenStudy (unklerhaukus):

hi !

OpenStudy (aa123):

@UnkleRhaukus can you help me with somthing

OpenStudy (aa123):

heres the question? http://openstudy.com/study#/updates/53626c58e4b0a83d8e0efe7a

OpenStudy (unklerhaukus):

Maybe before you try and find the inverses of A, you should check if A is invertible, can you find the determinant of A?

OpenStudy (unklerhaukus):

If the determinant of a matrix is 0, then this matrix is singular; it has no inverse